A tall man with a red tie is riding up in an elevator. He realizes that he is standing on a scale, which reads 700N, while the elevator is at rest. When the elevator begins to move, the scale reads 476N. a) Determine the acceleration of the elevator. b) How long will it take to reach 10m/s?
1
Expert's answer
2019-12-10T09:54:45-0500

The weight of man in elevator


"P=m(g+a)=P_0(1+a\/g)"

(a) The acceleration of elevator


"a=g\times \frac{P-P_0}{P_0}"

"=10\:\rm m/s^2\times \frac{476\:\rm N-700\: N}{700\:\rm N}=-3.2\:\rm m/s^2"

(b) The time that needed man to reach 10 m/s


"t=\frac{\Delta v}{a}=\frac{10\:\rm m/s}{3.2\:\rm m/s^2}=3.125\: \rm s"